What number Am i?
1)I have 4 digits, and they are all odd.
2)The digit in the tenths place is 3.
3)Add 4 to the number in the tenths place and you will have the number in the hundredths place.
4)The 2-digit whole number is greater than 10 and less than 20.If 121 is divided by this 2-digit whole number, the quotient is also that number.

Answer by Edwin McCravy(20056)

What number Am i?
1)I have 4 digits, and they are all odd.

↏↏.↏↏

2)The digit in the tenths place is 3.

↏↏.3↏

3)Add 4 to the number in the tenths place 
and you will have the number in the hundredths place.

            3+4 = 7  so

↏↏.37

4)The 2-digit whole number is greater than 10 and less than 20.
If 121 is divided by this 2-digit whole number, the quotient is 
also that number.

Since we know that 11×11 = 121, we know that 121÷11 = 11

So the answer is

11.37

Edwin